OpenStudy (anonymous):

On the graph of the equation 2x - y = 10, what is the value of the y-intercept? A.-10 B.-5 C.5 D.10

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Im asking this i don't know lol

OpenStudy (skullpatrol):

Do you know how to transform an equation by addition?

OpenStudy (anonymous):

Maybe why?

OpenStudy (skullpatrol):

You will need it to answer this question that's why.

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (skullpatrol):

Yes, why?

OpenStudy (anonymous):

can u tell me plz I'm not doing good I'm math

OpenStudy (anonymous):

in

OpenStudy (skullpatrol):

First subtract 10 from both sides of the equation.

OpenStudy (anonymous):

ok i get -8

OpenStudy (skullpatrol):

Where do you get -8 from?

OpenStudy (anonymous):

2 - 10

OpenStudy (skullpatrol):

But the 2 has an "x" multiplying it, that's why they write it "2x"?

OpenStudy (skullpatrol):

2x - y = 10

OpenStudy (skullpatrol):

First subtract 10 from both sides of the equation: 2x - y = 10 2x - y = 10 2x - y - 10 = 10 -10

OpenStudy (anonymous):

ok

OpenStudy (skullpatrol):

Now what do you get?

OpenStudy (anonymous):

-20

OpenStudy (skullpatrol):

10 - 10 =?= -20

OpenStudy (anonymous):

-10-10=-20

OpenStudy (skullpatrol):

First subtract 10 from both sides of the equation: 2x - y = 10 2x - y = 10 2x - y - 10 = 10 -10